# Towards the Terminator Economy: Assessing Job Exposure to AI through LLMs

### v2

@@ 6 Conclusions
A human evaluation has been performed to assess the approach’s effectiveness, revealing that an average of 71-75% of the time evaluators agree with the score and explanations provided by our indexes, whilst only 2% of the time tend to disagree fully. To ensure full reproducibility, our methodology relies only on top-performing open LLMs, and all occupation and task lists are accessible through an online demo. This would allow the community to track future assessments of improvements in new LLM models while integrating our findings to help individuals and organizations shape policies for AI integration into careers and the workforce.
Finally, while O*NET offers a comprehensive and widely recognized taxonomy of the labour market, its static nature may not capture emerging tasks as they emerge in the real-labour market. To address this limitation, our approach is designed to be flexible and adaptable to any user-defined taxonomy, enabling organizations to assess the impact of AI within their operational contexts accurately.
Online Demo. We have made the results of our framework shown in this paper available at https://terminatoreconomy.com/home.

### v3

@@ 6 Conclusions and Limitations
A human evaluation has been performed to assess the approach’s effectiveness, revealing that an average of 71-75% of time evaluators agree with the score and explanations provided by our indexes, whilst only 2% of time tend to disagree fully. To ensure full reproducibility, our methodology relies only on top-performing open LLMs, and all occupation and task lists are accessible through an online demo. This would allow the community to track future assessments of improvements in new LLM models while integrating our findings to help individuals and organizations shape policies for AI integration into careers and the workforce.
Finally, while O*NET offers a comprehensive and widely recognized taxonomy of the labor market, its static nature may not capture emerging tasks as they emerge in the real-labor market. To address this limitation, our approach is designed to be flexible and adaptable to any user-defined taxonomy, enabling organizations to assess the impact of AI within their operational contexts accurately.
Appendix A Appendix
A.1 Evaluating Consensus Among LLMs
To construct a single indicator, we took a conservative approach by assigning to each task the value of the rating with the highest frequency among the three models; if the three ratings differed, we chose the lowest. To assess the agreement between the three rates expressed by the LLMs, for each single task, we compute a consensus metric TASTLE2007531.
This process allows the three rates to be combined into a single measure of agreement. A higher consensus value indicates greater agreement between the observed rating values expressed for each task by the three different LLMs. The equation (4) shows the consensus calculation in which $LV_{k}$ represents the observed rating value, $p_{k}$ its relative frequency, $\mu_{LV}$ represents the weighted average of the $LV$ ratings using $p_{k}$ probabilities as weights, and $d_{L}V$ represents the scale size of the ratings adopted. The logarithmic function computes the impact of the normalised difference between each rating and the weighted average, moderated by the $d_{L}V$ dimension. The calculation uses a repeated summation for each k-th rate expressed for each individual task. Analogously, to estimate the similarity between the motivations provided by the LLMs, we compute the centroid of semantic cosine similarity rahutomo2012semantic, between them. The embedding vectors for the centroid computation is obtained using an open source Transformer model: as for the LLMs, we chose the Transformer model to be used in accordance with the Massive Text Embedding Benchmark (MTEB) Leaderboard by muennighoff-etal-2023-mteb. [^13] https://huggingface.co/spaces/mteb/leaderboard. Having English-language motivations, the choice fell on the UAE-Large-V1 https://huggingface.co/WhereIsAI/UAE-Large-V1, which represented an excellent compromise between effectiveness and efficiency, given its small size.
In particular, since higher cosine similarity values reflect higher semantic similarities between the text of the LLM motivations, we expect a strong correlation between the consensus metric and cosine similarity. The heat map represented in Figure 8 shows that both the values of cosine similarity and the consensus metric are extremely high, with an average close to 0.9 in both cases.
On the one hand, this suggests coherence between LLM-generated rates and the associated motivations, and on the other hand it adds robustness to our conservative approach in choosing between different models.
Acknowledgements
This paper is partially supported by the research activity of an EU H2020 Project PILLARS — Pathways to Inclusive Labor Markets - Grant agreement 101004703 - https://www.h2020-pillars.eu/
Contribution Statement
E.C., F.M., and M.M conceived the study; F.M., M.M., and A.S. developed the approach from a computational perspective E.C. performed the analyses from an economic perspective; F.M. managed the acquisition of funds. All authors wrote the initial manuscript draft, reviewed, and approved the final version of the manuscript.